When Mr. Kim returned from his fishing trip, his students inquired as to the length of his prize catch. Mr. Kim answered, "The head measures 9 inches. The tail was as long as the head and half the body. The body was as long as the head and tail." How long was Mr. Kim's prized fish?
Answer by richwmiller(17219)

h=9
t=h+b/2
b=h+t
f=h+b+t
f=72 inches that's 6 feet